# Hybrid RAG-empowered Multi-modal LLM for Secure Healthcare Data Management: A Diffusion-based Contract Theory Approach

### v1

@@ III-B Hybrid RAG-empowered MLLMs Agents for Data Management
Data management tasks in hospitals and the health center include data storage, analysis, and retrieval. When multi-modal healthcare data is gathered into the subchains, the MLLM agent categorizes the data by type and stores it in the appropriate databases. As healthcare data is needed, the MLLM agent retrieves and analyzes the data, ensuring it meets the specific requirements of the task[16]. To further enhance the ability of MLLMs to analyze multi-modal healthcare data, we design a hybrid multi-modal RAG module[18], which is integrated with a data sharing mechanism inspired by contract theory, ensuring that the MLLM data analysis is conducted with high quality and strong privacy protection, allowing secure and effective handling of multi-modal healthcare data. As shown in the right of Fig. 1, the workflow of the Hybrid RAG-empowered MLLMs Agents for Data Management is presented as follows:
@@
Step 1 (Store multi-modal healthcare data): Hospitals and the healthcare center collect available multi-modal healthcare data, convert them into vectors specific to each modality using an embedding model, and store these vectors in the local knowledge database with SQL tools[9].
@@
Step 2 (Retrieve multi-modal healthcare data): When a task query is received, the hybrid multi-modal RAG system uses the same embedding model from Step 1 to convert the query into a vector. Then it calculates similarity scores between the task query vector and the vectors within the knowledge database, retrieving and prioritizing the top K vectors that most closely match the task query[10].
@@
Step 3 (Re-rank the retrieved information): When all relevant information, particularly multi-modal data, is fed directly into the MLLMs, information overload can result and attention to key details can be reduced due to the inclusion of irrelevant content[10]. To address this, the system further screens the results by applying our Multi-modal Information Similarity (MIS) metric, which is calculated by
@@
where $f_{i}(\cdot)$ represents the similarity measure function between the task query and the source data in the database and is determined freely according to the requirements of the specific task. $x_{1}$ and $x_{2}$ are the unimodal data corresponding to the task query and the source data in the database, respectively, and the weight factor $w_{i}$ is used to characterize the proportion of each the similarity measure function $f_{i}(\cdot)$. When the results are re-ranked and filtered by MIS, the retrieved optimized healthcare information is then used to expand the context in the prompt.
Step 4 (Optimize the multi-modal inputs): Upon completing the retrieval process, we employ prompt engineering to optimize and synthesize a coherent prompt that integrates the original multi-modal task query with the retrieved multi-modal healthcare data[10, 26]. This refined prompt is then used as the input for the MLLMs, enhancing its ability to generate accurate and contextually relevant responses.
Step 5 (Generate the corresponding content based on the inputs): Upon receiving the multi-modal inputs, MLLMs connect each modal input to its respective pre-trained encoder model, and an adapter module is used to unify all processed embedding[20]. Then, the MLLMs use the pre-trained LLM to generate the corresponding content based on the inputs[10].
In the generation of MLLMs, hybrid RAG can enhance the generation quality by effectively incorporating relevant information from various sources. However, RAG does not improve the generalization ability of MLLMs, meaning that the model’s ability to apply learned information to new, unseen contexts remains limited, limiting its overall learning ability. To improve the output quality of MLLMs, an essential way is to continuously incorporate new healthcare data for training. Thus, we propose an incentive mechanism to encourage hospitals to share updated healthcare data in Section IV.
@@ IV Problem Formulation
In this section, we begin by developing a metric for healthcare data quality, followed by the formulation of utility functions for both healthcare data holders and the MLLM service provider. Finally, we propose a contract theory model to motivate healthcare data holders to contribute high-quality healthcare data.
The training of MLLMs relies heavily on a large volume of high-quality data[8]. Unfortunately, most healthcare data are stored in hospital databases in various regions. Without data sharing, these valuable resources remain untapped, hindering MLLM development. Furthermore, the effectiveness of MLLMs is directly influenced by the quality of the data used in training. Therefore, it is critical to implement an incentive mechanism that encourages hospitals to share healthcare data. Referring to [13], we consider the healthcare industry comprising hospitals in diverse regions and a health center as an example. The health center acts as the MLLM service provider, and the hospitals in diverse regions serve as the healthcare data holders, represented by a set of =$\mathcal{M}=\{1,\ldots,m,\ldots,M\}$. Initially, we propose a healthcare data quality metric through the AoI metric to assess the quality of healthcare data utilized for fine-tuning MLLMs. Subsequently, acting as the data task publisher, the MLLM service provider employs a contract theory model to encourage $M$ healthcare data holders to engage in data sharing [23].
@@ IV-A Healthcare Data Quality Metrics
To generalize, we define the size of healthcare data as $\ell$ (bytes) and the transmission rate between the health center and hospitals as $\tau$ (bytes per second). Hence, the transmission time of the healthcare data is $t_{trans}=\ell/\tau$ [23]. Meanwhile, we denote $t_{u}$ as the time of completing a consensus process among blockchains[13]. Therefore, we represent the length of a single time slot $t$ as $t=t_{trans}+t_{u}$ [13]. To maintain data freshness, each healthcare data provider $m$ periodically updates its healthcare data, with $\theta_{m}$ indicating the length of a single time slot in each update cycle. The refreshment of healthcare data happens in the initial time slot of the cycle. Referring to [28], the AoI for a data request made in the $i$-th time slot is $(i+1)t$ for $i=2,\ldots,\theta_{m}-1$, and for requests initiated in the first or last time slot, the AoI is $2t$. Due to the Poisson process[23, 13], data requests are equally likely to occur in any time slot, with a probability of $1/\theta_{m}$. Therefore, the average AoI for data sharing by healthcare data holder $m$ is given by
@@
where $\overline{A}_{max}$ represents the maximum permissible value for the AoI. The average AoI plays a critical role in the quality of MLLM services. Given that (2) is a convex function in relation to the update cycle $\theta_{m}$, increasing $\theta_{m}$ results in a decrease in AoI[23]. Consequently, there is a tradeoff in managing AoI, which can be optimized by modifying the update cycle.
@@ IV-B Healthcare Data Holder Utility
In the context of healthcare data sharing for MLLM services, the utility for each healthcare data holder $m$ is the difference between the reward $R_{m}$ and the cost $C_{m}$ incurred by data sharing tasks, expressed as $U_{m}=R_{m}-C_{m}$ [13]. According to [29], the cost for healthcare data holder $m$ is defined as $C_{m}=\xi_{m}f_{m}$ [23], with $f_{m}=\frac{1}{\theta_{m}}$ representing the update frequency and $\xi_{m}$ denoting the cost of each update [13]. Therefore, the utility of healthcare data holder $m$ is
@@
The MLLM service provider lacks precise knowledge of each healthcare data holder’s update cost because of the information asymmetry. To address this, the MLLM service provider classifies data holders into discrete types by using statistical distributions derived from historical data, and its expected utility will be optimized[13]. By classifying $M$ healthcare data holders into various types, we denote the $k$-th type healthcare data holder as $\delta_{k}=1/\xi_{k}$ and group them into a set $\mathcal{K}=\{\delta_{k}:1\leq n\leq K\}$, where a smaller update cost corresponds to a higher healthcare data holder type, and the healthcare data holder types are organized as $\delta_{1}\leq\delta_{2}\leq\cdots\leq\delta_{K}$. Thus, the utility of the type-$k$ healthcare data holder is given by
@@
IV-C MLLM Service Provider Utility
Due to the quality of MLLMs’ output being affected by the freshness of the healthcare data, large AoI leads to poor output for MLLMs and reduces the satisfaction of the MLLM service provider. Referring to [25], The satisfaction function for the MLLM service provider, based on type-$k$ healthcare data holders, is defined as
$\begin{split}S_{k}=\alpha\log(G(\overline{A}_{k})+1),\end{split}$ (6)
where $\alpha$ is the overall zero-shot accuracy of MLLMs for various services. For example, the zero-shot accuracy of LLaVA-Med across different domains is presented in Table II. Here, the value of $\alpha$ is determined by past experience when applied to various services[30].
@@ IV-D Contract Formulation
In this scenario, the MLLM service provider takes the lead in designing a set of contract items and offers them to $K$ healthcare data holders. Based on its type, each healthcare data holder selects the most appropriate contract item, denoted by $\Psi_{k}=\{(f_{k},R_{k}),k\in\mathcal{K}\}$, where $f_{k}$ represents the update frequency for the type-$k$ healthcare data holder, and $R_{k}$ is the reward given to the type-$k$ healthcare data holder as an incentive for its contribution. To guarantee that each healthcare data holder opts for the most advantageous contract item for its type, the designed contract must adhere to both Incentive Compatibility (IC) and Individual Rationality (IR) constraints.
@@
Traditional mathematical solutions may not be able to effectively adapt to the dynamic changes in the environment during data sharing. Consequently, we turn to GDM as a more efficient approach to determine the optimal contract[31], which we discuss in the following section.
@@ V-A1 State space
To find the optimal contract item, i.e., $(f_{k,}^{*},R_{k}^{*})$, $\begin{matrix}k&\in&\mathcal{K}\end{matrix}$, the system first adds Gaussian noise to the initial contract sample. In the current diffusion round $t(t=1,2,\ldots,T)$, the state space affecting the optimal contract design is defined as
@@ V-A2 Action space
As the MLLM service provider designs the contract items $\Psi$ to incentivize healthcare data holders to provide high-quality healthcare data, the action $\bm{a}^{t}$ at round $t$ is defined as
@@ V-A3 Immediate reward
15             Update the contract quality network by minimizing the loss function: $\mathcal{L}(\omega)=\frac{1}{N}\sum_{i=1}^{N}\big{(}r_{i}+\gamma q_{\varphi^{\prime}}^{\prime}(\bm{s}_{i},\bm{a}_{i}^{0})-q_{\varphi}(\bm{s}_{i},\bm{a}_{i})\big{)}^{2}$.
16             Update the contract generation network by computing the policy gradient: $\nabla_{\omega}\bm{\epsilon}_{\omega}\approx\frac{1}{N}\sum_{i}\nabla_{\bm{a}^{0}}q_{\varphi}(\bm{s},\bm{a}^{0})|_{\bm{s}=\bm{s}_{i}}\nabla_{\omega}\bm{\epsilon}_{\omega}|\bm{s}_{i}$.
@@
22return The trained contract generation network $\bm{\epsilon}_{\omega}$.
@@ V-B GDMs for Optimal Contract Design
Compared to DRL algorithms that directly optimize model parameters, GDMs enhance contract design through an iterative process of denoising the initial distribution[31]. The diffusion model network maps the environmental state to the contract design, which constitutes the contract design policy represented as $\bm{\pi}_{\omega}(\bm{a}|\bm{s})$ with parameters $\omega$. The contract design policy $\bm{\pi}_{\omega}(\bm{a}|\bm{s})$ designed to generate an optimal contract over multiple time steps can be expressed as[31]
@@
$\bm{a}^{t-1}|\bm{a}^{t}=\frac{\bm{a}^{t}}{\sqrt{\chi_{t}}}-\frac{\delta_{t}}{\sqrt{\chi_{t}(1-\bar{\chi_{t}})}}\bm{\epsilon}_{\omega}(\bm{a}^{t},\digamma,t)+\sqrt{\delta_{t}}\bm{\epsilon},$ (17)
where $\digamma$ is the conditional information during the denoising process. Referring to [31], we effectively train the contract design policy $\bm{\pi}_{\omega}$ to enhance the training quality of the contract generation network $\bm{\epsilon}_{\omega}$. Additionally, inspired by the concept of the Q-function [32], we introduce a contract quality network $q_{\varphi}(\bm{s},\Psi)$. The training of the contract quality network utilizes the double Q-learning technique to minimize the Bellman operator, involving two critic networks $q_{\varphi_{1}},q_{\varphi_{2}}$ and the corresponding target critic networks $q_{\varphi_{1}^{\prime}},q_{\varphi_{2}^{\prime}}$. We define $q_{\varphi}=\min\{q_{\varphi_{1}},q_{\varphi_{2}}\}$, and the optimal contract design policy that maximizes the expected cumulative utility of the client is expressed as
$\bm{\pi}=-\mathbb{E}_{\bm{a}^{0}\sim\bm{\pi}_{\omega}}[q_{\varphi}(\bm{s},\bm{a}^{0})].$ (18)
We define the target policy as $\bm{\pi}_{\omega^{\prime}}$, and the optimization of $\varphi_{i}$ for $i=1,2$ is performed by minimizing the following objective function:
$\begin{split}\mathbb{E}_{\bm{a}^{0}\sim\bm{\pi}_{\omega^{\prime}}}\left[\left\|\left(r(\bm{s},\bm{a}_{z})+\gamma\min_{i=1,2}q_{\varphi_{i}^{\prime}}(\bm{s},\bm{a}_{z+1}^{0})\right)-q_{\varphi_{i}}(\bm{s},\bm{a}_{z})\right\|^{2}\right],\end{split}$ (19)
where $\gamma$ represents the discount factor and $\bm{a}_{z}$ represents the action in the training step $z$.
The pseudo-code of the proposed GDM-based contract generation scheme is shown in Algorithm 1, which consists of three phases, and its computational complexity is $\mathcal{O}(|\omega|+|\varphi|+E_{max}Z_{max}(T|\omega|+|\varphi|))$. In the proposed GDM-based contract generation scheme, denoising techniques are employed to generate optimal contract designs [31]. By integrating exploration noise into the contract design and executing it, the process accumulates exploration experience, contributing to the enhancement of contract quality.
Figure 2: The examples of hybrid RAG-empowered medical MLLMs. In the proposed hybrid RAG-empowered medical MLLM, the RAG initially retrieves healthcare data using unimodal methods. Next, we re-rank the information using metrics such as structural similarity index measure[33], normalized cross-correlation, and bert score[34]. The detailed information is then combined with the task query and input into the MLLM to generate results.

### v2

@@ III-B Hybrid RAG-empowered MLLM Framework for Data Management in IoMT
Data management tasks in hospitals and the health center include data storage, analysis, and retrieval. When multi-modal healthcare data is gathered into the subchains, MLLMs categorize the data by type and store it in the appropriate databases. As healthcare data is needed, MLLMs retrieve and analyze the data, thus meeting the specific requirements of data management tasks[22]. To further enhance the ability of MLLMs to analyze multi-modal healthcare data, we design a hybrid multi-modal RAG module[18], which is integrated with a data sharing mechanism inspired by contract theory, ensuring that the MLLM data analysis is conducted with high quality and strong privacy protection, allowing secure and effective handling of multi-modal healthcare data. As shown in the right of Fig. 1, the workflow of hybrid RAG-empowered MLLMs for data management in IoMT is presented as follows:
@@
Step 1. Store multi-modal healthcare data: Hospitals and the health center collect available multi-modal healthcare data, convert them into vectors specific to each modality using an embedding model, and store these vectors in the local knowledge database with Structured Query Language (SQL) tools[12].
@@
Step 2. Retrieve multi-modal healthcare data: When a task query is received, the hybrid multi-modal RAG system uses the same embedding model from Step 1 to convert the query into a vector. Then, the system calculates similarity scores between the task query vector and vectors within the knowledge database, retrieving and prioritizing the top $K$ vectors that most closely match the task query[13].
@@
Step 3. Re-rank the retrieved information: When all relevant information, particularly multi-modal data, is fed directly into MLLMs, it can lead to information overload and may reduce attention to critical details due to the inclusion of irrelevant content[13]. To address this, the system further screens the results by applying our Multi-modal Information Similarity (MIS) metric, which is calculated by
@@
where $f_{i}(\cdot)$ represents the similarity measure function between the task query and the source data in the database and is determined freely according to the requirements of the specific task, $x_{1}$ and $x_{2}$ are the unimodal data corresponding to the task query and the source data in the database, respectively, and the weight factor $w_{i}$ is used to characterize the proportion of each the similarity measure function $f_{i}(\cdot)$. When the results are re-ranked and filtered by MIS, the retrieved optimized healthcare information is then used to expand the context in the prompt.
Step 4. Optimize the multi-modal inputs: Upon completing the retrieval process, we employ prompt engineering based on zero-shot prompting technology to optimize and synthesize a coherent prompt that integrates the original multi-modal task query with the retrieved healthcare data[13, 31]. By using ‘probability’ as a control keyword, this refined prompt enhances the credibility of the entire prompt and improves the MLLM’s ability to generate accurate and contextually relevant responses.
Step 5. Generate the corresponding content based on the inputs: Upon receiving the multi-modal inputs, MLLMs connect each modal input to its respective pre-trained encoder model, where a pre-trained linear projection adapter is employed to unify all processed embeddings[20, 32]. This linear projection adapter, trained on 600K image-text pairs from PMC-15M, standardizes the embeddings, allowing pre-trained LLMs to generate the corresponding content based on the inputs[32, 13].
In the generation of MLLMs, hybrid RAG enhances the generation quality by effectively incorporating relevant information from various sources. However, RAG does not improve the generalization ability of MLLMs, indicating that the ability of MLLMs to apply learned information to new and unseen contexts remains limited, limiting its overall learning ability. To improve the output quality of MLLMs, an essential way is to continuously incorporate new healthcare data for training. Thus, we propose an incentive mechanism to encourage data holders to share updated healthcare data.
@@ IV Problem Formulation
In this section, we begin by developing a metric for healthcare data quality, followed by the formulation of utility functions for both healthcare data holders and the MLLM service provider. Finally, we formulate a contract theory model to motivate healthcare data holders to contribute high-quality healthcare data.
The training of MLLMs relies heavily on a large volume of high-quality data[11]. Unfortunately, most healthcare data are stored in hospital databases in various regions. Without data sharing, these valuable resources remain untapped, hindering MLLM development. Furthermore, the effectiveness of MLLMs is directly influenced by the quality of the data used in training. Therefore, it is critical to implement an incentive mechanism that encourages hospitals to share healthcare data. Referring to [16], we consider there are multiple hospitals in diverse regions and a health center as an example. The health center acts as the MLLM service provider, and the hospitals in diverse regions serve as the healthcare data holders, represented by a set of =$\mathcal{M}=\{1,\ldots,m,\ldots,M\}$. Initially, we propose a healthcare data quality metric through the AoI metric to assess the quality of healthcare data utilized for fine-tuning MLLMs. Subsequently, acting as the data task publisher, the MLLM service provider employs a contract theory model to encourage $M$ healthcare data holders to engage in data sharing [26].
@@ IV-A Healthcare Data Quality Metrics
To generalize, we define the size of healthcare data as $\ell$ (bytes) and the transmission rate between the health center and hospitals as $\tau$ (bytes per second). Hence, the transmission time of healthcare data is $t_{trans}=\ell/\tau$ [26]. Meanwhile, we denote $t_{u}$ as the time of completing a consensus process among blockchains[16]. Therefore, we represent the length of a single time slot $t$ as $t=t_{trans}+t_{u}$ [16]. To maintain data freshness, each healthcare data provider $m$ periodically updates its healthcare data, with $\theta_{m}$ indicating the length of a single time slot in each update cycle. The refreshment of healthcare data happens in the initial time slot of the cycle. Referring to [34], the AoI for a data request made in the $i$-th time slot is $(i+1)t$ for $i=2,\ldots,\theta_{m}-1$, and for requests initiated in the first or last time slot, the AoI is $2t$. Due to the Poisson process[26, 16], data requests are equally likely to occur in any time slot, with a probability of $1/\theta_{m}$. Therefore, the average AoI for data sharing by healthcare data holder $m$ is given by
@@
where $\overline{A}_{max}$ represents the maximum permissible value for the AoI. The healthcare data quality metric plays a critical role in the quality of MLLM services. Given that (2) is a convex function in relation to the update cycle $\theta_{m}$, increasing $\theta_{m}$ results in a decrease in AoI[26]. Consequently, there is a tradeoff in managing AoI, which can be optimized by modifying the update cycle.
@@ IV-B Healthcare Data Holder Utility
In the context of healthcare data sharing for MLLM services, the utility for each healthcare data holder $m$ is the difference between the reward $R_{m}$ and the cost $C_{m}$ incurred by data sharing tasks, expressed as $U_{m}=R_{m}-C_{m}$ [16]. According to [35], the cost for healthcare data holder $m$ is defined as $C_{m}=\xi_{m}f_{m}$ [26], with $f_{m}=\frac{1}{\theta_{m}}$ representing the update frequency and $\xi_{m}$ denoting the cost of each update [16]. Thus, the utility of healthcare data holder $m$ is
@@
Due to information asymmetry, the MLLM service provider lacks precise knowledge of the update cost of each healthcare data holder. To address this, the MLLM service provider classifies data holders into discrete types by using statistical distributions derived from historical data, and its expected utility will be optimized[16]. By classifying $M$ healthcare data holders into various types, we denote the $k$-th type healthcare data holder as $\delta_{k}=1/\xi_{k}$ and group them into a set $\mathcal{K}=\{\delta_{k}:1\leq n\leq K\}$, where a smaller update cost corresponds to a higher healthcare data holder type, and the healthcare data holder types are organized as $\delta_{1}\leq\delta_{2}\leq\cdots\leq\delta_{K}$. Thus, the utility of the type-$k$ healthcare data holder is given by
@@
IV-C MLLM Service Provider Utility
Due to the quality of MLLMs’ output being affected by healthcare data freshness, large AoI leads to poor output for MLLMs and reduces the satisfaction of the MLLM service provider. Referring to [29], The satisfaction function for the MLLM service provider, based on type-$k$ healthcare data holders, is defined as
$\begin{split}S_{k}=\alpha\log(G(\overline{A}_{k})+1),\end{split}$ (6)
where $\alpha$ is the overall zero-shot accuracy of MLLMs for various services. For example, the zero-shot accuracy of LLaVA-Med as a medical LLM across different domains is presented in Table II. Here, the value of $\alpha$ is determined by past experience when applied to various services[32].
@@ IV-D Contract Formulation
In this scenario, the MLLM service provider takes the lead in designing a set of contract items and offers them to $K$ healthcare data holders. Based on its type, each healthcare data holder selects the most appropriate contract item, denoted by $\Psi_{k}=\{(f_{k},R_{k}),k\in\mathcal{K}\}$, where $f_{k}$ represents the update frequency for type-$k$ healthcare data holders, and $R_{k}$ is the reward given to type-$k$ healthcare data holders as an incentive for its contribution. To guarantee that each healthcare data holder opts for the most advantageous contract item for its type, the designed contract must adhere to both Incentive Compatibility (IC) and Individual Rationality (IR) constraints.
@@
Traditional mathematical solutions often struggle to effectively adapt to the complexity and dynamic changes inherent in data sharing environments[36]. In response, we leverage GDMs, a key component of GenAI, which excel not only in image generation but also in optimizing network performance[37, 6]. Building on similar approaches[38, 28], we employ GDMs as a more efficient solution for identifying optimal contracts. This approach capitalizes on the generative capabilities of GDMs to capture uncertainties and fluctuations in network conditions, allowing for more accurate identification of optimal contracts in real-time scenarios and effectively addressing the high-dimensional and intricate nature of the problem[17].
@@ V-A1 State space
To find the optimal contract item, i.e., $(f_{k,}^{*},R_{k}^{*})$, $\begin{matrix}k&\in&\mathcal{K}\end{matrix}$, the system first adds Gaussian noise to the initial contract sample. In the current diffusion round $t=1,2,\ldots,T$, the state space affecting the optimal contract design is defined as
@@ V-A2 Action space
As the MLLM service provider designs a contract $\Psi$ to motivate healthcare data holders to provide high-quality healthcare data, the action $\bm{a}^{t}$ at round $t$ is defined as
@@ V-A3 Immediate reward
14             Update the contract quality network by minimizing (19).
15             Update the contract generation network by computing the policy gradient (18).
@@
21return The trained contract generation network $\bm{\epsilon}_{\omega}$.
@@ V-B GDMs for Optimal Contract Design
Compared with DRL algorithms that directly optimize model parameters[39], GDMs can enhance contract design through an iterative process of denoising the initial distribution[38, 28]. The diffusion model network maps the environmental state to contract design, which constitutes the contract design policy represented as $\bm{\pi}_{\omega}(\bm{a}|\bm{s})$ with parameters $\omega$. The policy $\bm{\pi}_{\omega}(\bm{a}|\bm{s})$ designed to generate an optimal contract over multiple time steps can be expressed as
@@
$\bm{a}^{t-1}|\bm{a}^{t}=\frac{\bm{a}^{t}}{\sqrt{\chi_{t}}}-\frac{\delta_{t}}{\sqrt{\chi_{t}(1-\bar{\chi_{t}})}}\bm{\epsilon}_{\omega}(\bm{a}^{t},\bm{s},t)+\sqrt{\delta_{t}}\bm{\epsilon}.$ (17)
Referring to [38, 28], we effectively train the contract design policy $\bm{\pi}_{\omega}$ to enhance the training quality of the contract generation network $\bm{\epsilon}_{\omega}$. Additionally, inspired by the concept of the Q-function [40], we introduce a contract quality network $q_{\varphi}(\bm{s},\Psi)$. The training of the contract quality network utilizes the double Q-learning technique to minimize the Bellman operator, involving two critic networks $q_{\varphi_{1}},q_{\varphi_{2}}$ and the corresponding target critic networks $q_{\varphi_{1}^{\prime}},q_{\varphi_{2}^{\prime}}$. We define $q_{\varphi}=\min\{q_{\varphi_{1}},q_{\varphi_{2}}\}$, and the optimal contract design policy that maximizes the expected cumulative utility of the client is expressed as[28]
$\bm{\pi}=\arg\max_{\bm{\pi}_{{\omega}}}\mathbb{E}\Bigg{[}\sum_{z=0}^{Z}\gamma^{z}(r(\bm{s}_{z},\bm{a}_{z})-\varsigma\bm{\pi}_{\omega}(\bm{s}_{z})\log\bm{\pi}_{\omega}(\bm{s}_{z}))\Bigg{]},$ (18)
where $\gamma$ represents the discount factor, $\bm{a}_{z}$ represents the action in the training step $z$, and $\varsigma$ represents the temperature coefficient controlling the strength of the entropy.
We define the target policy as $\bm{\pi}_{\omega^{\prime}}$, and the optimization of $\varphi_{i}$ for $i=1,2$ is performed by minimizing the following objective function[28]:
$\begin{split}&\mathbb{E}_{(\bm{s}_{z},\bm{a}_{z},\bm{s}_{z+1},r_{z})\sim\mathcal{B}_{z}}\Big{[}\sum_{l=1,2}(r(\bm{s}_{z},\bm{a}_{z})-q_{\varphi_{l}}(\bm{s}_{z},\bm{a}_{z})\\ &\quad\quad\quad+\gamma^{z}(1-d_{z+1})\bm{\pi}_{\omega^{\prime}}(\bm{s}_{z+1})q^{\prime}_{\varphi^{\prime}}(\bm{s}_{z+1}))^{2}\Big{]},\end{split}$ (19)
where $\mathcal{B}_{z}$ is a mini-batch of transitions sampled from the experience replay memory $\mathcal{D}$ in the training step $z$ and $d_{z+1}$ is a 00-$1$ variable denoting the terminated flag.
The pseudo-code of the proposed GDM-based contract generation scheme is shown in Algorithm 1, which consists of three phases, and its computational complexity is $\mathcal{O}(|\omega|+|\varphi|+E_{max}Z_{max}(T|\omega|+|\varphi|))$. In the proposed GDM-based contract generation scheme, denoising techniques are employed to generate optimal contract designs [38, 28]. By integrating exploration noise into the contract design and executing it, the process accumulates exploration experience, contributing to the enhancement of contract quality.
Figure 2: A real case study of hybrid RAG-empowered medical MLLMs. In the proposed hybrid RAG-empowered medical MLLM, the RAG initially retrieves healthcare data using unimodal methods. Next, we re-rank the information using metrics such as structural similarity index measure[41], normalized cross-correlation, and bert score[42]. The detailed information is then combined with the task query and input into the MLLM to generate results.